OpenStudy (samsan9):

What is the value of the fourth term in a geometric sequence for which a1 = 15 and r = 1/3? Express your answer as a fraction.

OpenStudy (cwrw238):

nth term = a1 r^(n-1) so for example the third term = a1r^2 ehere a1 = first term and r = common ratio

OpenStudy (samsan9):

i got 5/9 is that right?

OpenStudy (cwrw238):

4th term = 15(1/3)^3 = 15 * 1/3^3 = 15/27 = 5/9

OpenStudy (cwrw238):

yes